# Image captioning in different languages

### v2

@@ 1.2 Additions
Table 1: Image description datasets available in languages other than English, with an indication of their source, and whether the descriptions were Translated or Independently collected. Asterisks indicate that the data is a subset of the original dataset. Flickr8K is the predecessor of Flickr30K, see Hodosh et al. 2013. For some recources multiple references are given since it is not clear which is the ‘main’ reference.
@@ 2 The list
Table 1 provides a list of non-English captioning datasets. [^1] The Crossmodal-3600 dataset (Thapliyal et al., 2022, 36 languages) is not listed in the table because it would result in too much repetition., [^2] Some authors opt not to translate the dataset but instead use an English-language captioning model and then translate the results (e.g. Tien et al. 2020). These works are also not in the table. Some observations:
@@
  • Although Table 1 differentiates between translated and independently collected descriptions, in practice this distinction may be blurred by annotators providing manual corrections of automatic translations, sometimes replacing the translations altogether.

### v3

@@ 1.2 Additions
Table 1: Image description datasets available in languages other than English (A-S), with an indication of their source, and whether the descriptions were Translated or Independently collected. Asterisks indicate that the data is a subset of the original dataset. Flickr8K is the predecessor of Flickr30K, see Hodosh et al. 2013. For some recources multiple references are given since it is not clear which is the ‘main’ reference. In the case of Assamese, Choudhury et al. (2023) manually corrected translation errors from the datasets provided by Nath et al. (2022), resulting in a revised (but highly similar) dataset.
Language Source T I Citation
@@
Table 2: Image description datasets available in languages other than English (T-Z), with an indication of their source, and whether the descriptions were Translated or Independently collected. Asterisks indicate that the data is a subset of the original dataset. Flickr8K is the predecessor of Flickr30K, see Hodosh et al. 2013. For some recources multiple references are given since it is not clear which is the ‘main’ reference.
@@ 2 The list
Tables 1 and 2 provide a list of non-English captioning datasets. [^1] The Crossmodal-3600 dataset (Thapliyal et al., 2022, 36 languages) is not listed in the table because it would result in too much repetition., [^2] Some authors opt not to translate the dataset but instead use an English-language captioning model and then translate the results (e.g. Tien et al. 2020). These works are also not in the table. Some observations:
@@
  • Although Tables 1 and 2 differentiate between translated and independently collected descriptions, in practice this distinction may be blurred by annotators providing manual corrections of automatic translations, sometimes replacing the translations altogether.